What is the fair market value of a bond with the following features: (a) par value of $1000, (b) coupon payment of $60, and (c) 10-year maturity? Assume that current interest rates are 5%.

a. $930.00.

b. $1077.21.

c. $1110.00.

d. $1000.00.
Answer #1

1- (1+r) Valueof Bond = Coupon * - MaturityValue (1+r)

Where r is the discounting rate of period 5%

And n is the no of periods­­­­­­­ 10 years

1- (1+0.05) 10 Valueof Bond = 60* — 1000 (1 + 0.05)10 0.0577

= 463.30 + 613.91

= $ 1077.21  

OPtion B is correct.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.